OpenStudy (anonymous):

Rework problem 22 from section 3.5 of your text involving inspection at a widget factory. Assume that the box contains 4 defective widgets and 16 acceptable widgets. There are 3 inspectors, each of whom selects a widget from the box at random, inspects it, and then replaces it in the box. Each inspector makes his or her inspection at a different time. What is the probability that at least one inspector will find a defective widget? Please help!

OpenStudy (anonymous):

If there are 20 in all, and there are only 3 selections, how many include at least one defective widget?

OpenStudy (anonymous):

I'm sorry?

OpenStudy (anonymous):

I really don't understand what I'm supposed to do here

OpenStudy (anonymous):

There is a 1 in 4 chance for one selection to have a defective widget

OpenStudy (anonymous):

Basicly\[\frac{ 4 }{ 16 }\] simplifies to \[\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

So\[\frac{ 1 }{ 4 }\times3\] Shows the total chance that one is defective

OpenStudy (anonymous):

i tried 3/4 already

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }=0.75\]

OpenStudy (anonymous):

I tried inputting that answer too

OpenStudy (anonymous):

the system says it's incorrect

OpenStudy (anonymous):

Hold on...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[4\div20=0.2\]\[0.2\times3=0.6\] So there is a 0.6 chance that one will be defective.

OpenStudy (anonymous):

Sorry about that.

OpenStudy (anonymous):

that's not right either :c

OpenStudy (anonymous):

One second, doing some reading.

OpenStudy (anonymous):

Is it 0.2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 }\] or\[\frac{ 4 }{ 20 }\]

OpenStudy (anonymous):

If that isn't the answer, than try bumping the question so someone else can see it, and hopefully help you with it.